Hi agian, I have asked this question on a number of occasions but have so far only received a small amount of help. Can someone please show me how to simplify the following? I need to give answers in the form a+bi ( a, b are real numbers)
a) i^15
b) (2-i) / (5+i)
Thanks, -Nick.

Answer by Alan3354(21584)

a) i^15
i^15 = i^3 = -1*i
= -i
------
b) (2-i) / (5+i)
--------
Multiply NUM and DEN by the conjugate of the DEN
= (2-i)*(5-i)/(5+i)*(5-i)
= (10 - 1 - 7i)/(25+1)
= (9-7i)/26